OpenStudy (anonymous):

if the point P(-3/5,y) lies on the unit circle and P is in the third quadrant, what does y equal?

OpenStudy (amistre64):

the unit circle is made of cos, sin coords

OpenStudy (anonymous):

Since the x value is negative, we are talking left of the y-axis (Q2 or 3). We are told Q# so that means that our y-coord must also be negative.  Consider than creating a triangle by drawing a radius and drawing a line up to the x-axis

OpenStudy (amistre64):

sin^2 + cos^2 = 1 ; was my first impression :) which is identical to MrYs instructions

OpenStudy (anonymous):

-4/5

OpenStudy (amistre64):

3-4-5 WAS a shortcut , cool lol

OpenStudy (anonymous):

Then, from the triangle example (and amistre) you can use pythagorean theorem to find the unknown side (which is also the y-coord) \[y^2=1^2+(\frac{-3}{4})^2\] Solving this gives you 4/5, but you have to remember that since it is in Q3, it has to be negative.